OpenStudy (anonymous):

Simplify: (–7) • (8) – (4) • (–2) a.58 b.16 c.–30 d.–48

OpenStudy (projects):

What do you believe your answer is?

OpenStudy (anonymous):

For the numbers in parentheses, or numbers next to each other, you multiply them both. (−7)(8) − (4)(−2), =−56 − (4)(−2), =−56 − (−8), What do you think the answer is?

OpenStudy (projects):

^^^^

OpenStudy (anonymous):

Then just do normal subtraction, also negatives cancel out subtraction signs. So its adding the (-8) as normal 8.

OpenStudy (anonymous):

So we have now, -56 + 8, Positives added to negatives make the negative number go down normally.

OpenStudy (anonymous):

-56 + 8, Using the information we gave you, the answer should be 8 down from -56, Making the solution, -48 I hope this helped!

OpenStudy (anonymous):

Thanks for giving me a medal, Have a nice day.